- 2/28-3/15: Health check and maintenance for VIS
PRM: One flag was dropped. -> leave it
- 3/4-14 PRM: Misalignment does not reach where beam hits damper.
3/6 PR3: range -> incline by 20-30mrad at mirror because of bistable.
3/15: BS: Yaw offset adjustment by hand. -> 3/5-6
3/12: BS ok
3/5 afternoon or 3/6 morning: Coolers for ITMs will be stopped for 2-3 hours to change water.
- 3/18-19: Re-alignment
- 3/19: Closing
- 3/20: pump down
1/25-2/20: Yend Pcal finalization
3/5: EYA remaining works by Enomoto, Inoue.
2/25-3/5: Yend WAB finalization
3/6: Y end closing (no laser hazard at least in the morning)
3/7: Y end pump down
3/7,8: leak test
2/11-2/14: Xend NAB installation
2/21: Removing FC, Payload release
2/27: ETMX release
2/23-3/6: Xend Pcal finalization 1
3/6: Beam position by IR and Green
3/5 AOS: EY and EX(if they have time) work.
3/6: Green X alignment (no IR), and no hazard at Y end.
- 3/11, 18,19: Xend WAB finalization.
3/11~3/12: ETMX tuning.
- 3/19: Heatlink etc.
- 3/19: ETMX finalization by Takahashi
- 3/18,19: ETMX alignement
- 3/21 closing
- 3/22 Pumping down
- Leak test
- 3/18~ : Night work starts, daily meeting will be moved to 1PM
- Summary of work will last for a while in the morning meeting.
- 2/18-2/22: PRMI commissioning
- 3/25~?: DRMI commissioning
- 3/25-4/19: Y arm commissioning
- 3/25-4/19: FPMI commissioning
- Night work does not start from 3/11 due to the 2weeks delay by VIS troubles.
- It does not mean that night work is prohibited but it means that officially night shift responsible person won't be assigned and we have no daily meeting at 1pm.
- Official night work starts 3/18 or 3/25, but not determined yet.
Report of work
- Type B: SR2: We found ghost beam of oplev for SR2, and we need several works to fix it and others.
- Type Bp: PRM: We found nothing was touching except for cable, and we found loosen screws. We will fix them today.
- Type A: ETMX will be fixed for
- MIF: Green laser was swapped. We noticed that the AC power was low, and that is the cause of mulfunction.
- Beat note is somehow low.
- OMC rack is being implemented specially for circuits.
- Green light injection for SR2 table by changing laser source will be done today. Aluminum foil will be put at SR2 table.
- CRY: Temperature of BF of ITMX was fixed and shows correct one. ITMY was connected but still show some fake one.
- Touching around Cryopayload will be checked today.